Circular No. 6624

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 G1 (MONTANI) Further precise CCD positions have been reported as follows:

1997 UT α_{2000} δ_{2000} m_{1} Observer
Apr. 11.13696 7^{h}17^{m}20.^{s}38 +19^{o}57′35′′.5 19.0 Montani
11.14571 7 17 20.71 +19 57 34.5 19.2  _{”}
11.17508 7 17 21.80 +19 57 31.0 19.1 Balam
11.17792 7 17 21.86 +19 57 31.1 19.0  _{”}
12.22310 7 18 00.11 +19 56 20.7 18.8  _{”}
12.22553 7 18 00.26 +19 56 20.4 18.8  _{”}
12.23037 7 18 00.43 +19 56 19.7 18.8  _{”}

J. Montani (Lunar and Planetary Laboratory). 0.9-m Spacewatch tele-

scope. Measurer J. V. Scotti. D. D. Balam (University of Victoria). 1.8-m Plaskett telescope.

The following parabolic orbital elements, from 13 observations Apr. 9–12, are extremely uncertain. Nevertheless, it seems likely that the comet is quite far from the earth and perhaps of short period.

T = 1997 Sept. 14.987 TT  ω = 241.^{o}457
Ω = 263.385 }2000.0
q = 4.30578 AU  i = 3.667
1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Apr. 12 7^{h}17.^{m}87 +19^{o}56.^{′}6 4.450 4.494 86.^{o}1 12.^{o}9 18.8
22 7 24.59 +19 43.5 4.573 4.472 77.9 12.7 18.8
May 2 7 32.39 +19 26.8 4.692 4.450 70.1 12.3 18.8
12 7 41.13 +19 06.3 4.805 4.430 62.5 11.7 18.9
22 7 50.66 +18 41.6 4.911 4.411 55.2 10.9 18.9
June 1 8 00.87 +18 12.7 5.007 4.394 48.0  9.9 18.9
11 8 11.62 +17 39.3 5.093 4.379 41.1  8.8 18.9

COMET 121P/SHOEMAKER-HOLT 2 Total visual magnitude estimates: Jan. 3.41 UT, 13.5 (A. Hale, Cloud-

croft, NM, 0.41-m reflector); 9.40, 13.6 (Hale); Feb. 2.90, 14.3 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 7.04, 13.7 (G. Bakos, Ráktanya, Hungary, 0.44-m reflector); 8.96, 13.9 (J. Carvajal, Madrid, Spain, 0.44m reflector); Mar. 1.82, 13.8: (Hornoch); 4.81, 14.1 (Hornoch); 7.98, 14.7 (K. Szentaskó, Ráktanya, Hungary, 0.44-m reflector); 12.93, 14.4 (Hornoch); Apr. 2.92, 13.7 (K. Sarneczky, Ráktanya, Hungary, 0.44-m reflector).

1997 April 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden